20. Once you balance the following redox reaction occurring in aqueous solution, there are how many water molecules on the right side? H2O2 + HClO → HCIOg + H2O a. 2 b. 4 c. 6 d. 8
Answer #1

reaction is
a H2O2 + b HClO -- > c HClO3 + d H2O

balancing H
2a + b = c + 2d (1)
balancing O
2a + b = 3c + d (2)
balancing Cl
b = c (3)
let b = 1
so, c = 1
put value of b and c in (1)
2a + 1 = 1 + 2d
a = d (4)
put value of b and c in (2)
2a + 1 = 3 + d
2a - 2 = d
from (4) a = d
2a - 2 = a
a = 2
so, d = 2

balanced equation is
2H2O2 + 1HClO -- > 1HClO3 + 2H2O
so, cofficient of water molecule = 2

Answer : 2